# Member, IEEE

### v2

@@ Abstract
The rapid detection of abnormal body temperatures in urban populations is essential for managing public health risks, especially during outbreaks of infectious diseases. Multi-drone thermal screening systems offer promising solutions for fast, large-scale, and non-intrusive human temperature monitoring. However, trajectory planning for multiple drones in complex urban environments poses significant challenges, including collision avoidance, coverage efficiency, and constrained flight environments. In this study, we propose an enhanced trust region sequential convex optimization (TR-SCO) algorithm for optimal trajectory planning of multiple drones performing thermal screening tasks. Our improved algorithm integrates a refined convex optimization formulation within a trust region framework, effectively balancing trajectory smoothness, obstacle avoidance, altitude constraints, and maximum screening coverage. Simulation results demonstrate that our approach significantly improves trajectory optimality and computational efficiency compared to conventional convex optimization methods. This research provides critical insights and practical contributions toward deploying efficient multi-drone systems for real-time thermal screening in urban areas. For reader who are interested in our research, we release our source code at https://github.com/Cherry0302/Enhanced-TR-SCO.
{IEEEkeywords}
@@ 2 Problem Formulation
2.1 System Model and Notations
@@
2.2 Objective Function Formulation
@@
The smoothness of the drone trajectories is crucial to avoid sudden and drastic maneuvers of the drones. We measure the smoothness of the trajectory of the $k$-th drone over the waypoints from $t=2$ to $t=T-1$ using the following formula:
@@
The efficiency of the flight paths is measured by minimizing the total flight distance of the drones. For the $k$-th drone, the total flight distance over the waypoints from $t=1$ to $t=T-1$ is calculated as:
@@ 2.3 Constraints Definition
2.3.1 Physical and Safety Constraints
@@
2.3.2 Multi-Drone Coordination Constraints
@@
2.3.3 Regulations and Performance Constraints
@@
3 Methodology
3.1 Introduction to the Trust Region Sequential Convex Optimization Algorithm
@@ 3.2.1 Convexification
where $\nabla_{x,y}h_{\text{obs}}$ denotes the gradient of the obstacle height function at the reference point. This linear approximation is valid within the trust region radius $\delta_{x,y}$, ensuring the constraint remains convex and tractable. The flight altitude constraint $z_{\text{min}}\leq z_{k}(t)\leq z_{\text{max}}$ and boundary constraints $10\leq x_{k}(t),y_{k}(t)\leq S-$ 1 are inherently linear and thus directly incorporated as convex constraints without modification.
@@
3.2.2 Adaptive Trust-Region Filtering for Constraint Reduction
@@
3.2.3 Higher-Order Soft Trust-Region Integration
@@
3.3 Mathematical Formulation and Optimization Approach
@@
4 Simulation and Results
4.1 Description of Simulation Environment and City Model
@@ 4.2 Results and Analysis
Figure 1 illustrates the optimized trajectories of five drones navigating a simulated urban environment designed for thermal screening. The environment, represented as a 50$\times$50 3D grid with 20-30% building density, imposes constraints such as a 20-50 m altitude range for effective infrared temperature detection and safe distances from obstacles. The trajectories demonstrate the enhanced TR-SCO algorithm’s ability to generate smooth, energy-efficient paths that avoid abrupt maneuvers, critical for maintaining stable drone orientation and accurate thermal imaging. Drones navigate around buildings and other obstacles while maintaining safe inter-drone spacing, showcasing the algorithm’s integration of convexified nonlinear constraints—like obstacle height approximations via Taylor expansion within trust regions—to ensure safety and feasibility. The paths prioritize coverage of pedestrian-dense areas, reflecting the algorithm’s dynamic waypoint allocation to minimize redundant screening and maximize efficiency in crowded zones, a key requirement for effective public health monitoring.
@@
Table 1: Performance Comparison for 5 Drones
@@
Table 1 presents a performance comparison of four trajectory planning methods for 5 drones in urban thermal screening, evaluating total path length, total coverage area, and calculation time. The enhanced TR-SCO algorithm achieves the shortest total path length at 859.11 meters, demonstrating superior energy efficiency by balancing obstacle avoidance, inter-drone spacing, and dynamic trust region adjustments to minimize detours. In contrast, the Original TR-SCO, MINLP, and SCS show longer paths due to less refined constraint handling: the original lacks terrain-adaptive safety margins, MINLP struggles with local minima in discrete optimization, and SCS suffers from unstable linearization without trust region bounds.
@@
Table 2 shows a proportional increase in total path length and coverage area, reflecting the need for more complex coordination. Despite the higher problem complexity, the enhanced TR-SCO algorithm maintains computational efficiency with a convergence time of 1.58 seconds, demonstrating scalability.
Table 2: The Enhanced TR-SCO Algorithm Performance for 5 and 10 Drones

### v3

# Enhanced Trust Region Sequential Convex Optimization for Multi-Drone Thermal Screening Trajectory Planning in Urban Environments

@@ Abstract
The rapid detection of abnormal body temperatures in urban populations is essential for managing public health risks, especially during outbreaks of infectious diseases. Multi-drone thermal screening systems offer promising solutions for fast, large-scale, and non-intrusive human temperature monitoring. However, trajectory planning for multiple drones in complex urban environments poses significant challenges, including collision avoidance, coverage efficiency, and constrained flight environments. In this study, we propose an enhanced trust region sequential convex optimization (TR-SCO) algorithm for optimal trajectory planning of multiple drones performing thermal screening tasks. Our improved algorithm integrates a refined convex optimization formulation within a trust region framework, effectively balancing trajectory smoothness, obstacle avoidance, altitude constraints, and maximum screening coverage. Simulation results demonstrate that our approach significantly improves trajectory optimality and computational efficiency compared to conventional convex optimization methods. This research provides critical insights and practical contributions toward deploying efficient multi-drone systems for real-time thermal screening in urban areas. For readers who are interested in our research, we release our source code at https://github.com/Cherry0302/Enhanced-TR-SCO.
Index Terms:
@@ II Problem Formulation
II-A System Model and Notations
@@
II-B Objective Function Formulation
@@
The smoothness of the drone trajectories is crucial to avoid sudden and drastic maneuvers of the drones. We measure the smoothness of the trajectory of the $k$-th drone over the waypoints from $t=2$ to t=T1t=T-1 using the following formula:
@@
The efficiency of the flight paths is measured by minimizing the total flight distance of the drones. For the $k$-th drone, the total flight distance over the waypoints from $t=1$ to t=T1t=T-1 is calculated as:
@@ II-C Constraints Definition
II-C1 Physical and Safety Constraints
@@
II-C2 Multi-Drone Coordination Constraints
@@
II-C3 Regulations and Performance Constraints
@@
III Methodology
III-A Introduction to the Trust Region Sequential Convex Optimization Algorithm
@@ III-B1 Convexification
where $\nabla_{x,y}h_{\text{obs}}$ denotes the gradient of the obstacle height function at the reference point. This linear approximation is valid within the trust region radius $\delta_{x,y}$, ensuring the constraint remains convex and tractable. The flight altitude constraint $z_{\text{min}}\leq z_{k}(t)\leq z_{\text{max}}$ and boundary constraints $0\leq x_{k}(t),y_{k}(t)\leq S-1$ are inherently linear and thus directly incorporated as convex constraints without modification.
@@
III-B2 Adaptive Trust-Region Filtering for Constraint Reduction
@@
III-B3 Higher-Order Soft Trust-Region Integration
@@
III-C Mathematical Formulation and Optimization Approach
@@
IV Simulation and Results
IV-A Description of Simulation Environment and City Model
@@ IV-B Results and Analysis
Figure 1 illustrates the optimized trajectories of five drones navigating a simulated urban environment designed for thermal screening. The environment, represented as a $\times50$ 3D grid with 20-30% building density, imposes constraints such as a 20-50 m altitude range for effective infrared temperature detection and safe distances from obstacles. The trajectories demonstrate the enhanced TR-SCO algorithm’s ability to generate smooth, energy-efficient paths that avoid abrupt maneuvers, critical for maintaining stable drone orientation and accurate thermal imaging. Drones navigate around buildings and other obstacles while maintaining safe inter-drone spacing, showcasing the algorithm’s integration of convexified nonlinear constraints—like obstacle height approximations via Taylor expansion within trust regions—to ensure safety and feasibility. The paths prioritize coverage of pedestrian-dense areas, reflecting the algorithm’s dynamic waypoint allocation to minimize redundant screening and maximize efficiency in crowded zones, a key requirement for effective public health monitoring.
@@
TABLE I: Performance Comparison for 5 Drones
@@
Table I presents a performance comparison of four trajectory planning methods for 5 drones in urban thermal screening, evaluating total path length, total coverage area, and calculation time. The enhanced TR-SCO algorithm achieves the shortest total path length at 859.11 meters, demonstrating superior energy efficiency by balancing obstacle avoidance, inter-drone spacing, and dynamic trust region adjustments to minimize detours. In contrast, the Original TR-SCO, MINLP, and SCS show longer paths due to less refined constraint handling: the original lacks terrain-adaptive safety margins, MINLP struggles with local minima in discrete optimization, and SCS suffers from unstable linearization without trust region bounds.
@@
Table II shows a proportional increase in total path length and coverage area, reflecting the need for more complex coordination. Despite the higher problem complexity, the enhanced TR-SCO algorithm maintains computational efficiency with a convergence time of 1.58 seconds, demonstrating scalability.
TABLE II: The Enhanced TR-SCO Algorithm Performance for 5 and 10 Drones